OpenStudy (anonymous):

Simplify (6/x)/(6/x+1)

OpenStudy (anonymous):

\[\frac{\frac{6}{x}}{\frac{6}{x+1}} = \frac{6}{x} * \frac{x + 1}{6} = \frac{x + 1}{x} = 1 + \frac{1}{x}\]